OpenStudy (anonymous):

ok I do not understand this algebra... 2w-7w=10 w=

OpenStudy (rational):

2w - 7w = 10 -5w = 10 w = -2

OpenStudy (anonymous):

Ok but I am trying to understand this problem and how to solve it... I am in Pre algebra and I can get some of it but some of it is so confusing to me

OpenStudy (rational):

think of "w" as cookies or some other object

OpenStudy (anonymous):

So I just subtract it then then divide it into the answer?

OpenStudy (rational):

Exactly !

OpenStudy (rational):

you can subtract like terms

OpenStudy (rational):

2w - 7w = -5w but you cannot subtract if they are not both same

OpenStudy (rational):

for example you cannot combine below : 2w - 7z

OpenStudy (anonymous):

oh ok so then there is no equation for 2w-7z?

OpenStudy (anonymous):

because they are not like terms correct?

OpenStudy (rational):

You're right 2w - 7z = 10 we cannot do anything here because they are not like terms

OpenStudy (anonymous):

Ok thank you so much for your help. I apreciate it and I am sure I will be back for more help... LOL

OpenStudy (rational):

you're welcome :)